### v1

# Faster shortest-path algorithms using the acyclic-connected tree

Abstract
This paper gives a fixed-parameter linear algorithm for the single-source shortest path problem (SSSP) on directed graphs. The parameter in question is the nesting width, a measure of the extent to which a graph can be represented as a nested collection of graphs. We present a novel directed graph decomposition called the acyclic-connected tree (A-C tree), which breaks the graph into a recursively nested sequence of strongly connected components in topological order. We prove that the A-C tree is optimal in the sense that its width—the size of the largest nested graph—is equal to the nesting width of the graph. We then provide a linear-time algorithm for constructing the A-C tree of any graph. Finally, we show how the A-C tree allows us to construct a simple variant of Dijkstra’s algorithm which achieves a time complexity of $O(e+n\log w)$, where $n$ ($e$) is the number of nodes (arcs) in the graph and $w$ is the nesting width. The idea is to apply the shortest path algorithm separately to each component in the order dictated by the A-C tree. We obtain an asymptotic improvement over Dijkstra’s algorithm: when $w=n$, our algorithm reduces to Dijkstra’s algorithm, but it is faster when $w\in o(n)$, and linear-time for classes of graphs with bounded width, such as directed acyclic graphs.
1 Introduction
1.1 Motivation
Finding the shortest paths from a node in a weighted graph $G$ is one of the oldest and most fundamental algorithmic research problems, whose modern applications range over everything from robotics [35] to route planning [6]. This problem, now called the Single-Source-Shortest-Path Problem (SSSP), originates with the work of Dijkstra [18], whose eponymous algorithm efficiently computes a shortest path tree to all nodes from a given fixed source node $s$ in $G$. Dijkstra’s algorithm (using Fibonacci heaps) runs in time $O(e+n\log n)$, where $n$ and $e$ are the number of nodes and arcs in the graph, and is known to be worst-case optimal for producing a list of nodes in sorted order of distance from the source node [21], a problem called Distance Ordering (DO). In a recent work [28], a new variant of Dijkstra’s algorithm using a ‘working-set’ heap was shown to be universally optimal for DO, a form of beyond-worst-case analysis which measures worst-case performance for each graph individually, up to reweighting of the arcs.
The SSSP problem is somewhat easier than the DO problem, because the nodes do not have to be sorted—Dijkstra’s algorithm also produces a shortest-path tree as well as the sorted order. However, the lower bounds for DO are derived from the problem of comparison sorting. By contrast, SSSP does not share this lower bound—instead, the complexity is only well-understood in special cases (see Section 1.3).
In fact, we know that classes of graphs exist where SSSP is strictly easier than DO. An important example is directed acyclic graphs (DAGs), where SSSP can be solved in linear time [53]. Inspired by [28], we would like to obtain new beyond-worst-case results for SSSP. That is, we want a shortest path algorithm that has the same time complexity as Dijkstra’s algorithm in the worst case, but outperforms Dijkstra’s when the given graph possesses additional structure. Our focus in this paper is on hierarchical structure in particular, that is, when a graph can be hierarchically decomposed into a nested collection of smaller graphs. This is an important property for applications, where the concept of modularity naturally leads to such a nested structure. We show later that nesting collectively generalises several natural graph ideas, including dominance, connected components and acyclicity (see Section 2). In particular, every DAG can be decomposed into a nested collection of graphs with two nodes each.
Example 1.1.
[…]
[width=0.8]figs/intro_ex3
[…]
[width=0.8]figs/intro_ex3_nested
Figure 1: A graph $G$, and a subset of its nodes (Fig. 1(a)). This set of nodes allows $G$ to be decomposed into a pair of graphs $G_{1}$ and $G_{2}$, with $G_{2}$ nested at a node $X$ of $G_{1}$ (Fig. 1(b)).
Consider the graph $G$ in Figure 1 with source node $s$. Here, the set of nodes $S=\{a,d,e,f,g,h\}$ has an interesting property: all arcs into $S$ go to $h$. This observation allows us to run Dijkstra’s algorithm more efficiently on this graph, by breaking it into a composition of two smaller graphs. We can treat this set as being a smaller single-source graph $G_{2}$ (with source $h$) nested within a node $X$ of an ‘outer’ graph $G_{1}$ (Figure 1(b)). Now we can run Dijkstra’s algorithm on $G_{1}$, and when we reach $X$, run Dijkstra’s algorithm recursively on $G_{2}$, beginning at its source $h$. The resultant time complexity is $O(e_{1}+v_{1}\log v_{1}+e_{2}+v_{2}\log v_{2})\leq O(e_{1}+e_{2}+(v_{1}+v_{2})\log\max(v_{1},v_{2}))$, where $v_{i}$ and $e_{i}$ are the number of nodes and arcs in $G_{i}$, respectively. By comparison, naïvely running Dijkstra’s algorithm on $G$ results in the larger running time of $O(e_{1}+e_{2}+(v_{1}+v_{2})\log(v_{1}+v_{2}))$. The complexity of Dijkstra’s algorithm is determined by the nonlinear heap operations—by splitting the problem into two subgraphs, we get two smaller heaps, therefore, the running time improves.
The key observation in Example 1.1 can be generalised to more than one layer of nesting. Indeed, deferring formal details to later sections, given a graph $G$ decomposed into a nested tree of subgraphs $G_{1},G_{2},\dots,G_{m}$, we can use the same recursive application of Dijkstra’s algorithm to compute the shortest paths in time $O((\sum_{i}e_{i})+(\sum_{i}v_{i})\log\max_{i}v_{i})$.
With this as motivation, we would like to able to algorithmically decompose a graph into a tree of smaller nested subgraphs. Generally, graphs can be decomposed in several distinct ways. We therefore seek a decomposition where the maximum size of the nested subgraphs is minimised, a quantity we call the ‘width’ $w$ of the decomposition. We arrive at the main goal of our paper: obtain beyond-worst-case improvements to SSSP by efficiently constructing a minimal-width decomposition given as a tree of nested graphs.
1.2 Contribution
In this paper, we present a linear-time algorithm for computing a minimal-width decomposition of a graph $G$ and use this decomposition to solve SSSP in time $O(e+n\log\operatorname{nw}(G))$ using a recursive form of Dijkstra’s algorithm [^1] Our construction allows for any SSSP algorithm to be used here. This means that all improvements on the overall problem (such as [28]), can be immediately extended to graphs of small nesting width.. Here, the parameter $\operatorname{nw}(G)$ is called the nesting width of $G$, defined as the minimal width among all nesting decompositions. Since $\operatorname{nw}(G)\leq n$, this SSSP algorithm is an asymptotic improvement over Dijkstra’s algorithm. In detail, our paper has four main contributions.
First, we formalise a hierarchical decomposition of a single-source graph, which we call a nested decomposition (Definition 2.7). We demonstrate its connection to modules (Definition 2.5), which are subsets of nodes which are reachable only through a single node (cf. Example 1.1). We define the width of a nested decomposition as the maximal size of the subgraphs in the decomposition, and define the nesting width $\operatorname{nw}(G)$ of a graph $G$ as the minimal width over all nested decompositions of $G$ (Definition 2.8).
Second, exploiting the connection with modules, we define a new decomposition on a single-source graph $G$ called the acyclic-connected tree (A-C tree) of $G$ (Definition 3.1). The nodes of the A-C tree are the nodes of $G$, with each node having one incoming arc and possibly several outgoing arcs to disjoint sets of nodes in the tree. See Figure 2(b) for an example. The descendants of a node $a$ are those nodes $b$ where all paths to $b$ contain $a$. The sets of nodes that are children of $a$ define strongly connected components of this subgraph, which are ordered in topological order. The name derives from the fact that this tree defines a set of recursively nested acyclic graphs of strongly connected components. We note that the A-C tree does not depend on the weights of $G$ and prove that it is an optimal decomposition in the sense that it has minimal width.
Third, to use the A-C tree for SSSP, we need to be able to construct it efficiently. Specifically, the construction must be asymptotically faster than $O(e+n\log n)$, which is the complexity of simply running Dijkstra’s algorithm. This is our third contribution: we give a linear-time algorithm, called A-C_tree, for computing the A-C tree. Our approach combines two well-established graph concepts with linear-time algorithms: strongly connected components [53] and dominator trees [4].
Fourth, combining the A-C tree with a recursive application of Dijkstra’s algorithm, we solve SSSP for any graph $G$ in time $O(e+n\log\operatorname{nw}(G))$. This is an asymptotic improvement of Dijkstra’s algorithm. Indeed, in the worst case, $\operatorname{nw}(G)=n\in O(n)$, and our time complexity $O(e+n\log\operatorname{nw}(G))$ coincides with Dijkstra’s algorithm. However, whenever $\operatorname{nw}(G)\in o(n)$, our time complexity is strictly smaller. In particular, if a class of graphs has bounded nesting width (like DAGs), we can solve the shortest path problem in linear time.
1.3 Related work
Constructing fast shortest path algorithms has been an active area of research since the ‘50s, with pioneering work given by [40, 49, 7, 20, 18]. In particular, [21] showed that Dijkstra’s algorithm [18] (with Fibonacci heaps) is worst-case optimal among all algorithms which visit the nodes in order of distance. If one instead only computes the shortest-path tree, then faster algorithms are possible, depending on assumptions. One approach is to constrain the weights, for example by assuming them to be integers [17, 55, 9, 12] or uniformly bounded [17, 2, 32, 33]. We differ from this line of work since our approach is entirely structural—we assume no restriction on the weights except being non-negative. Other work using structural assumptions include search algorithms for planar graphs [37, 30, 19, 34, 41], minor-free graphs [56], directed acyclic graphs (DAGs) [53], and most recently universal optimality [28]. Generalising the DAG case, [52] showed that SSSP can be solved in time $O(e+n\log k)$ where $k$ is the size of the largest connected component. We obtain this as a special case of our method, because the size of the largest connected component is always at least $\operatorname{nw}(G)-1$, and is often strictly larger. Finally, there are several recent algorithms using continuous optimisation tools [48, 42, 15, 8, 42, 57]. We work instead directly in a discrete setup to be closer to the original (discrete) problem.
This paper introduces the A-C tree as a hierarchical decomposition of a graph. Similar concepts include the low diameter decomposition (see [9, 12] and references within), which groups together strongly connected sets with respect to the weights (essentially, nodes that are close together should be grouped together), with probabilistic guarantees. In particular, [9, 12] also hierarchically extract acyclic parts using topological order between strongly connected sets to compute shortest path trees faster, but only for integral weights (possibly negative integers). Another line of research is the hierarchy-based approach developed by [43, 54, 44, 29]. This approach divides the graph into subgraphs splitting up the original SSSP into smaller SSSPs to search faster. The basic idea is to find subgraphs that are safe to search in greedily (ignoring the remaining graph) up to a certain distance. This idea is quite similar to how the A-C tree divides the SSSP into smaller problems. However, these hierarchy-type approaches ultimately rely on the weights, whereas our decomposition is entirely structural. Further, it is shown in [43] that any hierarchy-type algorithm must run in $\Omega(e+n\log n)$ time in the worst case on a certain family of directed graphs. It is straightforward to show that this family has bounded nesting width, and so our algorithm runs in linear time on these graphs.
The A-C tree proposed in this work uses the notion of a dominator tree, introduced by [45] and with the first algorithm given by [38]. Given a graph with a fixed source node $s$, a node $m$ is said to dominate $v$ if every path (from $s$) to $v$ goes through $m$ (Definition 2.2), and the dominator tree is the tree representing this partial order over all nodes (with root $s$). Efficient algorithms for computing the dominator tree include the classical near-linear time algorithm given by [36], and more sophisticated linear time algorithms [4, 13, 14, 25]. In our paper, the dominator tree essentially defines the ‘acyclic’ part of the A-C tree. In this way, the A-C tree can be seen as an extension of the dominator tree, where we also recursively extract strongly connected components. Dominator trees have been used in several applications, such as the design of efficient compilers [1, 16], modelling ecosystems [3], diagnostic test generation [5], 2-vertex connectivity tests and approximations [23, 24], influencing diffusion networks [26], finding strong bridges in graphs [31], diagnosing memory leaks [39], and computing reachability in graphs subject to constraints [46]. However, dominator trees have rarely been applied to shortest path problems, as we do in this paper. One exception is the recent work of [28] which we discussed in the introduction. There they use the dominator tree in a subroutine to simplify the graph as part of the constructing of a universally optimal algorithm for distance ordering in terms of number of comparison queries. Our usage is similar in flavour, though the dominator plays a much more central role in our definitions and results.
Our hierarchical approach using modules is inspired by the recent work of [11, 10]. In [10], the authors study the structure of finite state machines (FSMs) by developing an algorithm to decompose them into hierarchical finite state machines. They achieve this by identifying ‘modules’ in the FSM and using this to construct an equivalent nested hierarchy of machines. However, their work did not consider planning on FSMs. This was extended by [51, 50], who showed how to solve SSSP on weighted hierarchical FSMs. Similar to our paper, they exploit the modularity of the hierarchy to apply Dijkstra’s algorithm recursively. However, [51, 50] assumes a given hierarchical decomposition of machines, while in this paper we first efficiently construct such a decomposition.
There is a plethora of shortest path methods which rely on a preprocessing step to compute shortest paths faster at run-time [6]. However, the preprocessing step itself is typically computationally expensive. For example, the preprocessing step in [51, 50] is exponential in the number of machines in the hierarchy. Computing the A-C tree in this paper can conceptually be viewed as a preprocessing step, but with only a linear-time overhead.
Finally, the nesting width proposed in this paper is inspired by the use of treewidth [47] in the field of parameterised complexity. Treewidth is defined by ‘tree decompositions’, of which each graph can possibly have many. The width of any such decomposition is a bound for the complexity of many graph algorithms. The treewidth is the minimum width over all tree decompositions. While computing the treewidth is NP-hard, some classes of graphs are known to have bounded treewidth, allowing some otherwise NP-complete problems to be solved on them in polynomial time. In the same way, if a class of graphs has bounded nesting width, we can solve SSSP in linear time. In contrast to treewidth, computing the nesting width can be done very efficiently using the A-C tree.
1.4 Outline
The remainder of the paper is as follows. Section 2 presents the preliminaries. Section 3 describes the A-C tree. Section 4 proves that the A-C tree is a minimal-width decomposition. Section 5 presents an algorithm for computing the A-C tree and shows that it has linear-time complexity. This algorithm is then used in Section 6 to solve SSSP through recursive application of Dijkstra’s algorithm. An example illustrating the theory is given in Section 7. Finally, Section 8 concludes the paper.
2 Preliminaries
2.1 Graphs
In this paper, a graph refers to a weighted directed graph with a distinguished node called the source. Specifically, a graph is a tuple $G=(V,E,s)$ where $V$ is a set of nodes, $s\in V$ the source, and $E\subset V\times V$ a set of arcs. Most of our paper will use only this data, but to find shortest paths we will add a non-negative-valued weight function $\ell:E\to_{\geq 0}$ which stores the weight for each arc, with graph denoted by $G=(V,E,s,\ell)$.
We assume that all nodes in $G$ are reachable from $s$. This assumption is mild, because we could easily prune unreachable nodes prior to searching from $s$. Our goal is to solve the single-source shortest path problem.
Definition 2.1 (Single-source shortest path (SSSP)).
Given a graph $G=(V,E,s,\ell)$, construct the shortest path tree from $s$ to all other nodes in $V$.
We solve SSSP using a hierarchical approach. The idea is based on Bellman’s principle [7], which in this context means that if a node $c$ is on a shortest path $a\stackrel{{\scriptstyle}}{{\leadsto}}b$, then the subpaths $a\stackrel{{\scriptstyle}}{{\leadsto}}c$ and $c\stackrel{{\scriptstyle}}{{\leadsto}}b$ must also be shortest paths. Specifically, if we know that all paths from $a$ to $b$ contain $c$, we can reduce the problem of computing the shortest path $a\stackrel{{\scriptstyle}}{{\leadsto}}b$ to the problem of separately finding shortest paths $a\stackrel{{\scriptstyle}}{{\leadsto}}c$ and $c\stackrel{{\scriptstyle}}{{\leadsto}}b$. We formalise this using the dominance relation.
Definition 2.2.
[36] Let $G=(V,E,s)$ be a graph. If $a$ and $b$ are nodes, we say that $a$ dominates $b$ (written $a\geq_{s}b$) if every path $s\stackrel{{\scriptstyle}}{{\leadsto}}b$ contains $a$.
In the literature on control-flow graphs, this dominance relation is used to compute the dominator tree [38], since the relation defines a tree order on the nodes with root $s$.
Lemma 2.3 ([36]).
The relation $\geq_{s}$ is a tree-structured partial order on the nodes of $G$, rooted at $s$. That is, (1) $\geq_{s}$ is a partial order, (2) for all $a\in V$, $s\geq_{s}a$ and (3) if $a\geq_{s}c$ and $b\geq_{s}c$, then either $a\geq_{s}b$ or $b\geq_{s}a$.
Definition 2.4 (Dominator tree [36]).
The dominator tree of a graph $G=(V,E,s)$ is the transitive reduction of the ordering $\geq_{s}$.
The concept of dominance is closely related to—but not exactly the same as—the sets of nodes which can be treated as a recursively nested graph, as in Section 1.1. We formalise such sets through the notion of a module.
Definition 2.5 (Module).
A module of $G=(V,E,s)$ is a set $M\subseteq V$ of nodes such that there exists an $m\in M$ where for every arc $u\xrightarrow{}v$, if $u\not\in M$ and $v\in M$ then $v=m$. In other words, all arcs into $M$ go to $m$. We call $m$ the source of $M$  [^2] Note that ‘modules’ also have a distinct interpretation in graph theory [22, 27]. Here we use a form specific to single-source graphs, which is closer to the usage in [10, 11]..
We also use the convention that if a set $M\subseteq V$ contains the source $s$ of $G$, then $M$ is a module if and only if $M$ is a module with source $s$. Intuitively, one can think of an extra arc into $s$ from outside $G$, capturing that $s$ is the source of $G$.
Because we assume all nodes are reachable from $s$, every set that is not the whole node set $V$ must have at least one incoming arc. Also note that if $M$ is a module with source $m$, then $m\geq_{s}v$ for every $v\in M$. Every graph has at least some modules: the whole set $V$ is always trivially a module, as are the singleton sets $\{v\}$ for each $v\in V$. We call these the trivial modules.
Modules have some nice algebraic properties. In particular, the following will be useful.
Lemma 2.6.
If $M$ and $H$ are overlapping modules (they intersect but neither contains the other), then $M\cup H$ and $M\cap H$ are both modules.
Proof.
Let $M$ and $H$ be modules with sources $m$ and $h$ respectively. Let $x\in M\cap H$. We know $m\geq_{s}x$ and $h\geq_{s}x$ so either (1) $m\geq_{s}h$ or (2) $h\geq_{s}m$ by Lemma 2.3. Assume w.l.o.g that (1) holds. Then any path $s\stackrel{{\scriptstyle}}{{\leadsto}}x$ contains first $m$, then $h$, then $x$. Because the subpath $h\stackrel{{\scriptstyle}}{{\leadsto}}x$ does not contain $m$, but $x\in M$, we must have $h\in M\cap H$. Claim: if $m\in M\cap H$ and $h\in M\cap H$ then $m=h$. To prove the claim, suppose that $h\in M\cap H$ and $m\in M\cap H$. Since there are paths $s\stackrel{{\scriptstyle}}{{\leadsto}}m$ and $s\stackrel{{\scriptstyle}}{{\leadsto}}h$, there must be an arc $u\xrightarrow{}v$ with $v\in M\cup H$ and $u\not\in M\cup H$ [^3] Unless $s\in M\cup H$. However, if $s\in M\cup H$, then $m=s$ or $h=s$ and since both $m$ and $h$ are in $M\cap H$, $s\in M\cap H$. Thus, $s$ must be the source of both $M$ and $H$ and the claim follows.. We know that $v=m$ or $v=h$, so $v\in M\cap H$. But then because $M$ and $H$ are modules, we must have $v=m=h$, completing the claim. If follows from this claim that either $m=h$, or $m\in M\setminus H$ and $h\in M\cap H$. Now let $u\xrightarrow{}v$ be such that $u\not\in M\cap H$ and $v\in M\cap H$. If $u\not\in M$, then $v=m$ as $M$ is a module, and so $m\in M\cap H$ implies $m=h$. If instead we have $u\not\in H$, this implies that $v=h$ as $H$ is a module. Hence $M\cap H$ is a module with source $h$. Similarly, let $u\xrightarrow{}v$ be such that $u\not\in M\cup H$ and $v\in M\cup H$. If $v\in M$, then $v=m$ as $M$ is a module. If $v\in H$, then $v=h$ as $H$ is a module. However, $h\in M\cap H$, so in this case $v\in M$ which implies that $v=m=h$. Hence $M\cup H$ is a module with source $m$. ∎
Finally, as discussed in the introduction, SSSP can be solved in linear time for DAGs. A simple linear algorithm for this is to first compute the topological order of the nodes and then visit the nodes in this order, updating distances to successor nodes as we visit each node [53]. This subroutine will play a key role in our algorithms since we work with nested acyclic graphs.
2.2 Nesting width
In Example 1.1, we could treat the set $S=\{r,d,e,f,g\}$ as a nested graph $G_{2}$ with source node $r$ precisely because all arcs into $S$ goes to $r$. In other words, we could represent this as a nested graph precisely because $S$ is a module. Note that if $S$ were not a module, say because an arc $b\xrightarrow{}d$ existed in Figure 1(a), then we would not be able to assign a unique source node to $G_{2}$. Hence the sets of modules define nesting decompositions of a graph.
Definition 2.7 (Nesting decomposition).
Let $G=(V,E,s)$ be a graph. A nesting decomposition of $G$ is a collection of non-overlapping modules which includes all trivial modules (the whole set $V$, and the singletons $\{v\}$ for each $v\in V$).
Each module $M$ in a nesting decomposition $F$ has a unique maximal module partition, which consists of the modules $H\in F$ such that $H\subset M$ and there is no module $K\in F$ with $H\subset K\subset M$. This is a necessarily a partition because we include the trivial singleton modules, and it is unique because all modules in the set are non-overlapping. The width of a nesting decomposition is the largest size of the maximal module partition for any module in the collection.
Definition 2.8 (Nesting width).
The width of a given nesting decomposition is the maximum size of a maximal module partition of any module in the decomposition. The nesting width of a graph $G=(V,E,s)$, written $\operatorname{nw}(G)$, is the minimum width over all nesting decompositions of $G$.
A nesting decomposition $F$ of $G$ can be equivalently represented by a tree. Indeed, consider the graph whose nodes are the modules in the decomposition $F$, with an arc $M\xrightarrow{}H$ if $M\supset H$ and no $K\in F$ such that $M\supset K\supset H$. This graph defines a tree whose leaves are the singleton modules. In this form, the width of the decomposition is the maximum number of children of any node in this tree.
Note that every graph has at least one trivial nesting decomposition, consisting of only the trivial modules $V$ and $\{v\}$ for each $v\in V$. As a tree, this is a single root node with all singletons as its children. In particular, we always have the upper bound $\operatorname{nw}(G)\leq n$. Generally, though, graphs can have many distinct nesting decompositions. We want to find the ‘best’ nesting decomposition, whose width is minimal and thus equal to the nesting width.
3 The Acyclic-Connected Tree
Nesting decompositions are defined by modules, which are themselves related to the dominance relation $\geq_{s}$. This is the first ingredient of our decomposition: using the dominator tree to compute the relation $\geq_{s}$. However, dominance alone is not enough to construct a minimal-width nesting decomposition. Consider the following example: a graph with nodes $s,a_{i},b_{i}$ for $i$ from 1 to $N$, where there are arcs from $s$ to $a_{1}$ and $b_{1}$, and arcs from $a_{i}$ and $b_{i}$ to both $a_{i+1}$ and $b_{i+1}$. The dominator tree of this graph is the trivial one-layer tree with root $s$. However, this is acyclic, so a nesting decomposition of width $2$ exists (follows from Theorem 4.1). More generally, while the dominator tree tells us that all modules have source node $s$, this isn’t sufficient to identify them.
In this example we can construct a minimal-width decomposition by topological sorting of the graph, given it is acyclic. Our main theorem proves that these two ideas—dominators and topological sorting—are always sufficient to construct a minimal-width decomposition of any graph. Collectively, we will use these to define the acyclic-connected tree (A-C tree) of the graph.
Definition 3.1 (A-C tree).
Let $G=(V,E,s)$ be a graph, and $T$ be its dominator tree. For any node $a$ in $T$, let $D(a)$ denote its descendants and $C(a)$ its children. For each $a$, we define a graph $G_{a}$ whose node set is $C(a)$, with an arc $u\xrightarrow{}v$ in $G_{a}$ if $u\neq v$ and there is an arc from $D(u)$ to $D(v)$ in $G[D(a)]$ (the induced subgraph of $G$ consisting of descendants of $a$). The acyclic-connected tree (A-C tree) of $G$ is a map $a\mapsto S^{a}_{1},\dots,S^{a}_{k_{a}}$ taking each node $a$ to the sequence of strongly connected components of $G_{a}$ in topological order [^4] There may be many topological orders, so this is not unique. However, the choice of ordering will not affect any analysis, so we will treat it as unique for simplicity.. We depict this as in Figure 2.
[…]
[width=0.8]figs/intro_plain
[…]
[scale=.8]figs/AC_tree_new
Figure 2: The graph $G$ from Example 1.1 (Fig. 2(a)) and its associated A-C tree (Fig. 2(b)), where the children of a node are topologically ordered from left to right. The child of each node $x$ in the tree is a set of nodes which represent a strongly connected component of $G_{x}$. We represent this by grouping the nodes in a rectangle (e.g., $e$ and $d$). We omit this rectangle for clarity if the strongly connected component is a singleton.
In the next section, we prove that the A-C tree defines a minimal-width nesting decomposition of the graph, and show how the A-C tree can be computed in linear time.
4 A-C trees are minimal-width nesting decompositions
In this section we explain why the A-C tree gives us an minimum-width nesting decomposition of a graph. Formally, we have the following result.
Theorem 4.1.
Let $T_{AC}$ be the A-C tree of a graph $G=(V,E,s)$, so $T_{AC}$ maps each node $a$ to a sequence of sets $S^{a}_{1},\dots,S^{a}_{k_{a}}$. Then the family of sets $AC:=\{\{a\}\cup\bigcup_{i=1}^{j}D(S^{a}_{i}):a\in V,j\in\{1,\dots,k_{a}\}\}$ is a minimum-width nesting decomposition of $G$, where $D(S_{i}^{a})$ is the set of descendants of the set $S_{i}^{a}$ in the A-C tree.
Proof.
We begin by showing that $AC$ indeed defines a nesting decomposition, and its width is equal to $1+\max_{a,i}|S^{a}_{i}|$. To do this we must show that every set is a module, and that no two sets overlap.
Fix $a$, and let $S_{1}^{a},S_{2}^{a},\dots,S_{k_{a}}^{a}$ be the associated sequence of strongly connected components of $G_{a}$. The sets $\{a\}\cup\bigcup_{i=1}^{j}D(S^{a}_{i})$ for $j\in\{1,\dots,k_{a}\}$ form a nested increasing sequence, with $\{a\}\cup\bigcup_{i=1}^{j}D(S_{i}^{a})\subset\{a\}\cup\bigcup_{i=1}^{k}D(S_{i}^{a})$ for $j<k_{a}$. The singleton $\{a\}$ is a module, and all nodes in $D(S_{i}^{a})$ are dominated by $a$. Because the strongly connected components $S_{i}^{a}$ are in topological order with respect to $G_{a}$, any arcs into $D(S_{j}^{a})$ must originate in $\{a\}\cup\bigcup_{i=1}^{j-1}D(S_{i}^{a})$. Hence, by induction, each $\{a\}\cup\bigcup_{i=1}^{j}D(S_{i}^{a})$ is a module. This gives us a collection of non-overlapping sets for every $a$.
Finally suppose that a node $v$ is in two sets $M$ and $H$ in $AC$, which have distinct source nodes $a$ and $b$. Then $v$ is dominated by both $a$ and $b$, and so by Lemma 2.3 either $a\geq_{s}b$ or $b\geq_{s}a$. Without loss of generality, we assume the former. In particular, $b$ is a descendant of some child $S_{i}^{a}$ of $a$. But then all descendants of $b$ are also in $D(S_{i}^{a})$, and so we have that $H\subseteq M$. Hence $AC$ is a nesting decomposition of $G$.
Pick a module $M:=\{a\}\cup\bigcup_{i=1}^{j}D(S^{a}_{i})$ in $AC$. We can precisely identify the size of its maximal module partition. First, the set $\{a\}\cup\bigcup_{i=1}^{j-1}D(S^{a}_{i})$ is a module in $AC$, and a subset of $M$, and by construction it is maximal. Hence, the remaining modules in the maximal module partition must be subsets of $D(S_{j}^{a})$ with sources other than $a$. The nodes $v$ in $S_{j}^{a}$ are the children of $a$ in the dominator tree. For each such $v$, the set $D(v)=\{v\}\cup\bigcup_{i=1}^{k_{v}}D(S^{v}_{i})$ is a module in $AC$. Because $v$ is a child of $a$ in the dominator tree, $v$ cannot be dominated by any other elements of $S_{j}^{a}$. In particular this means that the module $D(v)$ must be maximal inside $M$. Hence the maximal module partition of $M$ in $AC$ is exactly
$\left\{\{a\}\cup\bigcup_{i=1}^{j-1}D(S^{a}_{i}),D(v_{1}),\dots,D(v_{m})\right\}$ (1)
where $v_{1},v_{2},\dots,v_{m}$ are the nodes in $S^{a}_{j}$. Hence, the size of the maximal module partition of $M$ is equal to $1+m=1+|S^{a}_{j}|$. The width of $AC$ is the size of the largest maximal module partition of any module in $AC$, which is $1+\max_{a,i}|S^{a}_{i}|$. We denote this quantity by $w_{AC}$.
We will now show that no other nesting decomposition can have strictly smaller width than $w_{AC}$. For contradiction, suppose that a nesting decomposition $T$ exists with width $w<w_{AC}$. Note that every decomposition has width at least two, so this implies that $w_{AC}$ is at least three.
Since $w_{AC}=1+\max_{a,i}|S^{a}_{i}|$, we can choose some $\alpha$ and $k$ such that $|S^{\alpha}_{k}|=w_{AC}-1$. Let $K$ be the smallest module in $T$ that contains all of $S^{\alpha}_{k}$. By assumption, the maximal module partition of $K$ in $T$ has size at most $w<w_{AC}$. Therefore, there must exist a module $M$ in the maximal module partition of $K$ that contains a subset of $S^{\alpha}_{k}$ but not the whole set. In particular, we have two nodes $v_{1}\in S^{\alpha}_{k}\setminus M$ and $v_{2}\in S^{\alpha}_{k}\cap M$. Now, because $S^{\alpha}_{k}$ is strongly connected (and the definition of $G_{\alpha}$), there is a path $v_{1}\stackrel{{\scriptstyle}}{{\leadsto}}v_{2}$ in $D(S^{\alpha}_{k})$. For $M$ to be a module, its source $\beta$ must be contained in $v_{1}\stackrel{{\scriptstyle}}{{\leadsto}}v_{2}$ (since we go from $v_{1}\notin M$ to $v_{2}\in M$). Moreover, by definition of the dominator tree, there is also a path $s\stackrel{{\scriptstyle}}{{\leadsto}}v_{2}$ that does not contain any of the nodes of $v_{1}\stackrel{{\scriptstyle}}{{\leadsto}}v_{2}$ except $v_{2}$, and since $\beta$ must also be contained in $s\stackrel{{\scriptstyle}}{{\leadsto}}v_{2}$, [^5] This follows since $s\notin K$ but $v_{2}\in K$. To see the former, note that if $s\in K$, then $\beta=s$. But $\beta$ is contained in the path $v_{1}\stackrel{{\scriptstyle}}{{\leadsto}}v_{2}$ in $D(S^{\alpha}_{k})$, so $s\in D(S^{\alpha}_{k})$, which is a contradiction since $|S^{\alpha}_{k}|\geq 2$. we get that $\beta=v_{2}$. Therefore, $M\subseteq D(v_{2})$. Since $M$ was arbitrary and the maximal module partition of $K$ must cover $S^{\alpha}_{k}$, we get that the number of modules in this partition that overlaps $S^{\alpha}_{k}$ must be at least $|S^{\alpha}_{k}|$ and these modules are all contained in $D(S^{\alpha}_{k})$. However, $K$ must also have nodes outside of $D(S^{\alpha}_{k})$. [^6] To see it, assume $K$ is contained in $D(S^{\alpha}_{k})$. Since $K$ contains $S^{\alpha}_{k}$, the source $\beta$ of $K$ must belong to $S^{\alpha}_{k}$. However, since $|S^{\alpha}_{k}|\geq 2$, we may pick another node $v\neq\beta$ in $S^{\alpha}_{k}$. By definition of the dominator tree, there exists a path $s\stackrel{{\scriptstyle}}{{\leadsto}}v$ that does not contain $\beta$, a contradiction since $v\in K$. Therefore, there exist at least one other module $M^{\prime}$ in the maximal module partition of $K$ that accounts for the nodes outside of $D(S^{\alpha}_{k})$. We conclude that the maximal module partition of $K$ must have size at least $|S^{\alpha}_{k}|+1=w_{AC}$, so $w\geq w_{AC}$, a contradiction. ∎
5 Computing the A-C tree
We proved in Theorem 4.1 that the A-C tree defines a nesting decomposition of minimal width. However, in order to use this to compute shortest paths we must also compute the A-C tree very efficiently. In this section we combine some classical linear-time algorithms (dominator trees, depth-first-search and strongly connected components) to achieve an algorithm which computes the A-C tree in linear time (Algorithm 1).
The first (and most technical) step of A-C tree computation is construction of the dominator tree. Fortunately, several decades of algorithmic work has established linear-time algorithms for this problem [4]. The second step, recalling Definition 3.1, is to group the children $C(a)$ of a node $a$ in the dominator tree into components, defined by reachability in the induced subgraph $D(a)$. The main challenge of this step is efficiently constructing the induced graphs $G_{a}$ whose nodes are $C(a)$ and where there is an arc $u\xrightarrow{}v$ in $G_{a}$ if there is a path $u\stackrel{{\scriptstyle}}{{\leadsto}}v$ in $G[D(a)]$. In Algorithm 2, we show how to construct the graphs $G_{a}$ for every $a$ in linear time using a variant of depth-first search. Once each graph $G_{a}$ has been constructed, we can run a linear-time strongly connected components algorithm (such as [53]) on each $G_{a}$ from $a$ to generate the components of $G_{a}$ in topological order, giving the children of $a$ in the A-C tree.
Input : Graph $G=(V,E,s)$
Output : A-C tree $T_{AC}$
1 Create an empty dictionary $T_{AC}$;
2 Construct the dominator tree $T$ of $G$;
3 $graphs\leftarrow\texttt{DominanceGraphs}{(G,s,T)}$;
4 for each node $a$ do
       $K_{1},K_{2},\dots,K_{m}\leftarrow SCCs(graphs[a])$;
        // Topologically sort $G_{a}$
5       Set $T_{AC}[a]\leftarrow K_{1},K_{2},\dots,K_{m}$;
6
7return $T_{AC}$;
8
Algorithm 1 A-C_tree
Input : Graph $G$, with source $s$ and dominator tree $T$
Output : $G_{p}$ for each node $p$ given by $graphs$
1 Run RecursiveDominanceGraphs ($s$);
2 $s[p]\leftarrow p.\textnormal{{graph}}$ graph for each node $p$;
3 return $graphs$;
4 RecursiveDominanceGraphs $(v)$
5       $v.\textnormal{{found}}\leftarrow True$;
6
7      for each arc $(v,w)$ in $T$ do
8             if $w.\textnormal{{found}}=False$ then
9                   $v.\textnormal{{current}}\leftarrow w$;
10                   RecursiveDominanceGraphs $(w)$
11      for each arc $(v,w)$ in $G\setminus T$ do
12             $p\leftarrow w.\textnormal{{parent}}$;
13             if $p.\textnormal{{current}}\neq w$  then
14                   add the arc $(p.\textnormal{{current}},w)$ to $p.\textnormal{{graph}}$;
15
Algorithm 2 DominanceGraphs
Theorem 5.1.
Given a graph $G=(V,E,s)$, Algorithm 1 constructs its A-C tree in linear time $O(n+e)$.
Proof.
Algorithm 1 begins by constructing the dominator tree $T$ of $G$, which can be done in linear time [4]. Then we run the subroutine DominanceGraphs$(G,s,T)$ (Algorithm 2).
(Claim: Algorithm 2 computes $G_{a}$ for every $a$.) Algorithm 2 performs a depth-first search (DFS) on the graph $G$ starting at $s$. However, the key property is that the specific DFS tree it follows is the dominator tree $T$. It is immediate that this algorithm is linear time, as it examines each node and arc exactly once.
[…]
Conversely, assume we add an arc $x\xrightarrow{}y$ to $a.graph$ on line 2, while examining some arc $v\xrightarrow{}y$. We want to show that $x\xrightarrow{}y$ is then an arc of $G_{a}$. Note that $y.parent=w.parent=a$, so $y\in C(a)$. Moreover, $x=a.current$ and thus, $x\in C(a)$. Since $x$ and $y$ are distinct by line 2 (meaning $v$ is not dominated by $y$), it remains to show that $v\in D(x)$. To see this, observe that $v\in D(a)$, since the arc $v\xrightarrow{}w$ goes to $w\in D(a)$ and $w\neq a$. Because $v$ is the current node in the DFS and $x$ is on the current DFS path from $a$, we get that $v\in D(x)$. Thus, $x\xrightarrow{}y$ is an arc of $G_{a}$.
We conclude that for each $a$, Algorithm 2 results in $a.graph$ being equal to $G_{a}$.
Finally, we iterate through each node $a$ and run Tarjan’s strongly connected components algorithm [53] on $graph[a]$, which we know is equal to $G_{a}$. This algorithm runs in linear time, and returns the strongly connected components $K_{1},K_{2},\dots,K_{m}$ of $G_{a}$ in topological order. Note that because of the tree structure, each node is topologically sorted exactly once, ensuring the overall procedure is still linear time. The resultant map $T_{AC}$ is the A-C tree, by Definition 3.1. ∎
6 Shortest paths from the A-C tree
As we showed in the introduction, the A-C tree allows us to construct a nesting decomposition of the graph, and from a nesting decomposition we can obtain a solution to SSSP by running a standard shortest path algorithm on each nested graph. This gives us an $O(e+n\log\operatorname{nw}(G))$ time algorithm. However, we can also achieve this running time in a much simpler and more direct way from the A-C tree, just by modifying a few lines of Dijkstra’s algorithm.
Our method, which we call Recursive Dijkstra’s, is described in Algorithm 3. The idea is simply that, upon visiting a node $a$, we first recursively search its children $K^{a}_{1},K^{a}_{2},\dots,K^{a}_{m}$, where again we search recursively when we reach new nodes. Each component $K^{a}_{i}$ contains its own priority queue, which are updated individually.
Theorem 6.1.
Recursive Dijkstra’s solves SSSP in $O(e+n\log\operatorname{nw}(G))$ time.
Input : Graph $G=(V,E,s)$
Output : Shortest distances and path tree given by dist and parent
1 Create two dictionaries dist and parent;
2 For each $v$, set $\textnormal{{parent}}[v]\leftarrow None$ and $\textnormal{{dist}}[v]=\infty$;
3 Set $\textnormal{{dist}}[s]\leftarrow 0$;
4 Compute $T_{AC}\leftarrow\texttt{A-C\_tree}(G)$;
5 Assign a priority queue to each component $K^{a}_{i}$ with the nodes in $K^{a}_{i}$;
6 Run RecursiveDijkstra $(s)$;
7 return dist, parent ;
8 RecursiveDijkstra $(v)$
9
10      $K_{1},K_{2},\dots,K_{m}\leftarrow T_{AC}[v]$;
11       for each $i$ from 1 to $m$ do
12             ComponentDijkstra $(K_{i})$;
13
14 ComponentDijkstra $(\kappa)$
15       $Q\leftarrow\kappa.\textnormal{{queue}}$;
16       while $Q$ is not empty do
17             $(dist,v)\leftarrow Q.extract\_min()$;
18             $\textnormal{{dist}}[v]\leftarrow dist$;
19             RecursiveDijkstra $(v)$;
20             for each successor $w$ of $v$ do
                   $L\leftarrow\textnormal{{dist}}[v]+\textnormal{{weight}}(v,w)$ ;
                    // Potentially new distance to $w$
21                   if $L<\textnormal{{dist}}[w]$ then
                         $\textnormal{{dist}}[w]\leftarrow L$ ;
                          // Update distance to $w$
                         $\textnormal{{parent}}[w]\leftarrow v$ ;
                          // Save new shortest path to $w$
22                         $K\leftarrow$ component $K^{a}_{i}$ containing $w$ in $T_{AC}$;
                         $K.\textnormal{{queue}}.decrease\_key(w,L)$ ;
                          // Update priority queue
23
Algorithm 3 Recursive Dijkstra’s
Proof.
Overall, Recursive Dijkstra’s runs very similarly to the standard setup of Dijkstra’s algorithm, with two key differences: (1) for each node $a$ in the A-C tree it searches each connected component $K^{a}_{i}$ separately in topological order, and (2) it applies this process recursively using the A-C tree structure. In Algorithm 3, we highlight the differences from Dijkstra’s algorithm in blue. We will explain these below.
First, we show that solving SSSP separately over strongly connected components works. Suppose the dominator tree is trivial, with $s$ dominating all nodes but no other dominance relations. Then this algorithm reduces to solving SSSP by running Dijkstra’s separately on each connected component $K_{1},K_{2},\dots,K_{n}$ in topological order. We will first show that this produces the same output as running Dijkstra’s algorithm on the entire graph collectively. The shortest paths to nodes in $K_{1}$ does not depend on any other $K_{i}$, because no paths exist from $K_{i}$ to $K_{1}$. Hence Dijkstra’s algorithm on $K_{1}$ alone successfully finds the shortest paths to all nodes in $K_{1}$. Then we run it on $K_{2}$, having previously found all shortest paths to $K_{1}$. But this operates the same as running Dijkstra’s algorithm on $K_{1}\cup K_{2}$, where we already have solved shortest paths to $K_{1}$. Again, the larger $K_{i}$ do not affect the shortest paths, so again this correctly identifies shortest paths to $K_{1}\cup K_{2}$. The remainder of the argument proceeds identically, by induction.
Now we argue that this works in the recursive setting as well. Suppose that we are searching from a node $a$ in $T_{AC}$. Assume by induction that until this point, running Recursive Dijkstra’s on any node $x$ correctly identifies the shortest paths from $x$ to all nodes in $D(x)$. We will show that this holds as well for $a$. Let $K^{a}_{1},K^{a}_{2},\dots,K^{a}_{m}$ be the components of $a$, which again we will search in topological order. The only difference from the base case is that an arc $x\xrightarrow{}y$ in $G_{a}$ does not necessarily correspond to an arc $x\xrightarrow{}y$ in $G$—instead, it might correspond to an arc $u\xrightarrow{}w$ where $u\in D(x)$ and $w\in D(y)$. Suppose this is the case, for some $x$ and $y$ with $y\in K^{a}_{i}$. First, recall that because $D(y)$ is a module, $w=y$. Then, by the inductive hypothesis, we have already found the shortest path from $x$ to $u$, because $u\in D(x)$. Hence we can update the distance to $y$ in $K^{a}_{i}.queue$ by using $dist[u]+weight(u,y)$ (line 3). This operation takes place while searching recursively from $u$. That is, once we have run Recursive Dijkstra’s on $x$, we have correctly updated the distances to all successors $y$ in $G_{a}$. Hence searching the components in topological order in $G_{a}$ works as before, allowing us to correctly compute shortest paths from $a$ to all nodes in $D(a)$.
Having proved the correctness, it remains to show the time complexity of Algorithm 3. To this end, note first that all subroutines in Algorithm 3 are done in linear time, except running Dijkstra’s algorithm (given by ComponentDijkstra) on each strongly connected component. We compute this time complexity. Totally analogous to Section 1.1, assuming $m$ strongly connected components $K_{1},\dots,K_{m}$ in the A-C tree, the time complexity to run Dijkstra’s algorithm on all these components is $O\big{(}(\sum_{i=1}^{m}e_{i})+(\sum_{i=1}^{m}n_{i})\log(\max_{i}n_{i})\big{)}$, where $n_{i}$ and $e_{i}$ is the number of nodes and edges in $K_{i}$. We note that $\sum_{i=1}^{m}n_{i}=n$ and $\sum_{i=1}^{m}e_{i}\leq e$, where the latter holds since some arcs may not be in the components. Moreover, $\max_{i}n_{i}=\max_{i}|K_{i}|=w_{AC}-1=\operatorname{nw}(G)-1\leq\operatorname{nw}(G)$ by the proof of Theorem 4.1. We conclude that the time complexity to run Dijkstra’s algorithm on all components is
$O\left(\left(\sum_{i=1}^{m}e_{i}\right)+\left(\sum_{i=1}^{m}n_{i}\right)\log\left(\max_{i}n_{i}\right)\right)\leq O\left(e+n\log\operatorname{nw}(G)\right).$
This is also the time complexity of Algorithm 3 since all other subroutines are executed in linear time. ∎
7 Example
Consider again Example 1.1 from the introduction. Running algorithm A-C_tree (Algorithm 1) gives us the A-C tree of $G$ in Figure 2(b). By Theorem 4.1, this A-C tree defines a minimal-width nesting decomposition of $G$ with width 3. This $G$ demonstrates that nesting decomposition can be quite non-trivial, even in strongly connected graphs. A small-width decomposition can be found because $G$ has an ‘almost acyclic’ structure.
The algorithm also identifies the nesting bottleneck in this structure, which is the strongly connected set $\{e,d\}$. Indeed, when computing shortest paths using Recursive Dijkstra’s (Algorithm 3), the maximum quueu length used in the algorithm is equal to $\max_{i}|K_{i}|=|\{e,d\}|=2$, where maximum is taken over all strongly connected components $K_{i}$ in the A-C tree. This is a significant improvement compared to naïvely running Dijkstra’s algorithm, with a single large queue of length $9$ containing all nodes.
8 Conclusion
8.1 Summary
In this paper, we introduced a graph decomposition called the A-C tree to provide parametrised complexity bounds for SSSP. The A-C tree breaks up a single-source graph $G$ into a recursively nested sequence of strongly connected components in topological order. We showed that the A-C tree is optimal in the sense that it achieves the nesting width $\operatorname{nw}(G)$, the minimal width of any nesting decomposition of $G$. We also presented a linear-time algorithm for constructing the A-C tree, which we used to design a SSSP algorithm called Recursive Dijkstra’a with time complexity $O(e+n\log(\operatorname{nw}(G)))$.
8.2 Future Work
Our results demonstrate that, even on this classical and fundamental problem, new graph ideas can still lead to new insight. Among other things, we highlight how despite the important progress on the complexity of the distance ordering problem (DO), the—arguably more natural—SSSP problem is open in many cases. The fact that a combination of simple linear-time graph algorithms can lead to non-trivial advances on this problem suggest that there is room for significant progress in future work. The direct follow-up questions are the following:
First, classifying the classes of graphs with bounded nesting width is important. In a sense, these graphs are ‘somewhat acyclic’ (generalising the DAG case), and as a result, SSSP can be solved in linear time. As the example in this paper illustrate, these classes of graphs are surprisingly rich, and are therefore worth classifying. Second, it would be interesting to understand the frequency that graphs with low nesting width occur in (sparse) random graphs. Thirdly, our approach is entirely structural and does not depend on the arc weights. This resembles universal optimality in the sense of [28]. Our results could be a first step towards building a universally optimal algorithm for SSSP, similar to the results of [28] for DO. Fourthly, and similarly, our independence from weights could be relevant in application domains where, for example, the weights can change over time but the graph topology does not. One motivating example is road networks, where traffic conditions change over time.
Acknowledgements
The authors would like to thank Mihalis Yannakakis and Christos Papadimitriou for valuable feedback on the paper.